What is the price of a $1,000 par value semi-annual bond with 12 years to maturity and a coupon rate of 3.3% and a yield-to-maturity of 8.6%?
Answer #1

Information provided:

Future Value= $1,000

Time= 12 years*2= 24 years

Coupon payment= 3.3% 0.033/2= 0.0165*1,000= $16.50

Yield to maturity= 8.6%/2= 4.3%

The question is solved with the help of a financial calculator by entering the below data:

FV= 1,000; N= 24; PMT= 16.50; I/Y= 4.30%

Press CPT and PV to calculate the price of the bond.

The price of the bond is $608.09.